OpenStudy (anonymous):

Solve the following system of equations. 3x - 2y = 6 6x - 4y = 14

OpenStudy (anonymous):

MUTIPLY first equation by 2

OpenStudy (ganpat):

something is wrong..

OpenStudy (anonymous):

the answer could be 'no solution' or 'Infinitely Many Solutions' that could be the problem

OpenStudy (anonymous):

6x-4y=12................(1) 6x-4y=14................(2) subtacting the two means 0=-2 that means any value of x and y are possible

OpenStudy (anonymous):

It has no solution.

OpenStudy (anonymous):

ya sorry i was to write that means no value of x and y is possilble which simultaneously satisfies both the equations

OpenStudy (anonymous):

thus,no solution is answer